# Fine-tuning Neural Network Quantum States

### v2

@@ Acknowledgements.
Figure 6: Relative energy error $\Delta\varepsilon$, compared to DMRG, plotted as a function of the optimization steps for the $J_{1}$-$J_{2}$ Heisenberg model [refer to Eq. (4)] with $J_{2}/J_{1}=0.6$ on a cluster of $N=100$ sites. The orange curve represents the variational energy obtained using a Restricted Boltzmann Machine (RBM) with $K=384$ hidden neurons. The blue curve depicts the same network with the addition of the Marshall Sign Rule as a prior for the sign structure. In contrast, the green curve is obtained by optimizing the same network on top of the hidden representation $\bm{z}$ generated by the Transformer with hyperparameters $h=12$, $d=192$, $n_{l}=4$ at $J_{2}/J_{1}=0.4$.
@@
We focus on the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)], with $N=100$ sites, fixing the value of the frustration ratio to $J_{2}/J_{1}=0.6$. We consider a Restricted Boltzmann Machine (RBM) with $K$ hidden neurons [see Eq. (2)]. This network is employed in two distinct manners: trained directly on the physical configurations $\bm{\sigma}$, and trained on the hidden representations $\bm{z}_{p}$, which are generated by a pretrained ViT at $J_{2}/J_{1}=0.4$. As depicted in Fig. 6, using the hidden representations (green curve) achieves an accuracy of $\Delta\varepsilon\approx 10^{-3}$, which is two orders of magnitude higher compared to the same network defined directly on the physical configurations ($\Delta\varepsilon\approx 10^{-1}$, orange curve). The difference primarily arises from the physical properties of the system that are encoded in the hidden representations, such as sign structure, amplitudes, and symmetries. Furthermore, given that the sign structure at $J_{2}/J_{1}=0.4$ is well approximated by the MSR, we optimize an RBM, directly on the physical configurations, but implementing the Marshall sign prior (blue curve). This RBM achieves an accuracy of $\Delta\varepsilon\approx 10^{-2}$, underscoring that the information compressed in the hidden representation exceeds that provided by the Marshall sign prior.
.4 Optimization protocol
During the pretraining stage, we perform $N_{\text{opt}}=10000$ optimization steps. Then, during the fine-tuning stage, the number of steps is reduced to $N_{\text{opt}}=3000$. For all the simulations, we estimate stochastically the observables choosing a number of samples of $M=3000$. The optimization of the variational parameters is performed with the Stochastic Reconfiguration (SR) method [54]. In particular, working with variational states featuring approximately $P=10^{6}$ parameters, we employ the alternative formulation of SR [21, 55] efficient in the regime $P\gg M$ (available in NetKet [56] under the name of VMC_SRt.). We use a cosine decay scheduler for the learning rate, setting the initial value to $\tau=0.03$.
.5 Variational energies
In Table 1, we report the ground state energies of the $J_{1}$-$J_{2}$ Heisenberg model in Eq. (4) on a cluster of $N=100$ sites with periodic boundary conditions for various frustration ratios $J_{2}/J_{1}$. They are obtained through three distinct methodologies: DMRG, ViT trained from scratch, and ViT pretrained at $J_{2}/J_{1}=0.4$ and subsequently fine-tuned for other frustration ratios. Notably, the fine-tuned ViT exhibits remarkable accuracy when compared to DMRG results, reaching an accuracy $\Delta\varepsilon\lesssim 10^{-3}$ for all the values of the frustration ratio in the interval $J_{2}/J_{1}\in[0.1,0.6]$.
$\bm{J_{2}/J_{1}}$
DMRG ViT Fine-tuning
0.10
-0.425417395
-0.4254174
-0.425218
0.20
-0.408572967
-0.4085728
-0.408453
0.30
-0.393126745
-0.3931204
-0.393059
0.40
-0.380387370
-0.3803726
-0.380370
0.60
-0.380804138
-0.3807913
-0.379902
Table 1: Variational ground state energies for the $J_{1}$-$J_{2}$ Heisenberg chain with system size $N=100$. The DMRG computations are conducted employing a bond dimension up to $\chi=10^{3}$ under periodic boundary conditions. For both instances involving ViT, one trained from scratch and another pretrained at $J_{2}/J_{1}=0.4$ followed by fine-tuning, the Monte Carlo error attributed to finite sampling effects affects the last digit of the reported results.

### v3

@@ Appendix B Memory Efficiency in Fine-tuning and Pretraining Processes
Figure 7: Relative error in energy $\Delta\varepsilon$, compared to DMRG, plotted as a function of the optimization steps for the $J_{1}$-$J_{2}$ Heisenberg model [refer to Eq. (4)] with $J_{2}/J_{1}=0.6$ on a system of $N=100$ sites. The orange curve represents the variational energy obtained using a RBM with $K=384$ hidden neurons and $77568$ parameters. The blue curve depicts the same network with the addition of the Marshall Sign Rule as a prior for the sign structure. In contrast, the green curve is obtained by optimizing the same network on top of the hidden representation $\bm{z}$ generated by the Transformer with hyperparameters $h=12$, $d=192$, $n_{l}=4$ at $J_{2}/J_{1}=0.4$. Inset: Relative error in energy $\Delta\varepsilon$ of a RBM trained with the MSR prior as a function of the number of parameters. For comparison, the accuracy of the finetuned network is also shown.
Appendix C Connecting the features to order parameters
In this appendix we examine the Ising model in a transverse field [see Eq. (3)] aiming to establish a connection between the features learned by the ViT optimized at $g/\Gamma=1$ and the magnetization order parameter that controls the phase transition. To achieve this, we consider a fixed batch of physical spin configurations $\{\bm{\sigma}_{i}\}\sim|\Psi(\bm{\sigma};\theta_{p})|^{2}$. We then compute the corresponding hidden representation and perform principal component analysis (PCA) on it. In Fig. 6, we plot the principal component against the local magnetizations of the spin configurations, i.e., $\sum_{i=1}^{N}\sigma_{i}$. The two quantities exhibit a strong correlation. This organization of configurations in the feature space simplifies the description of the physics of the system, allowing for an easy transition from the ordered phase to the disordered phase. It would be interesting to extend this analysis to other systems, where the order parameter is associated to an off-diagonal operator in the computational basis.
Appendix D Comparison with Marshall Sign Prior
@@
A relevant question is trying to understand which kind of prior information is encoded in the features generated by the pretrained network. We focus on the $J_{1}$-$J_{2}$ Heisenberg model on a chain [see Eq. (4)], with $N=100$ sites, fixing the value of the frustration ratio to $J_{2}/J_{1}=0.6$. We consider a Restricted Boltzmann Machine (RBM) with $K$ hidden neurons [see Eq. (2)]. This network is employed in two distinct manners: trained directly on the physical configurations $\bm{\sigma}$, and trained on the hidden representations $\bm{z}_{p}$, which are generated by a pretrained ViT at $J_{2}/J_{1}=0.4$. As depicted in Fig. 7, using the hidden representations (green curve) achieves an accuracy of $\Delta\varepsilon\approx 10^{-3}$, which is two orders of magnitude higher compared to the same network defined directly on the physical configurations ($\Delta\varepsilon\approx 10^{-1}$, orange curve). The difference primarily arises from the physical properties of the system that are encoded in the hidden representations, such as sign structure, amplitudes, and symmetries. Furthermore, given that the sign structure at $J_{2}/J_{1}=0.4$ is well approximated by the MSR, we optimize an RBM, directly on the physical configurations, but implementing the Marshall sign prior (blue curve). This RBM achieves an accuracy of $\Delta\varepsilon\approx 10^{-2}$, underscoring that the information compressed in the hidden representation exceeds that provided by the Marshall sign prior. Despite increasing the number of parameters in RBMs, their performance remains inferior to the fine-tuned network due to the poor scaling behavior of the relative error in energy with the growth of network parameters and complicated structure of the landscape with a lot of local minima emerging when increasing the number hidden neurons (refer to the inset of Fig. 7).